OpenStudy (anonymous):

. A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach? Show your work. Round answer to the nearest hundredths.

OpenStudy (chawkins01):

Just use the Pythagorean theorem (a^2+b^2=c^2 whereas a=3 and c=10)

OpenStudy (anonymous):

oh thanks

OpenStudy (chawkins01):

No problem